A quiet day on the ice. – University of Copenhagen

NEEM > Field diaries folder > UK diaries 2009 > A quiet day on the ice.

07 June 2009

A quiet day on the ice.

Lou and Sverrir mark out the position for a storage garage.

Lou and Sverrir mark out the position for a storage garage.

The morning was very quiet; but eventually people began to work. The whole science trench is now ready for action; but some problems with the electronics for the DEP and line scanner make us wait half a day.

Plastic bags, pencils, forms, saws and instructions are ready. People are ready. We begin slowly. People have to learn the different tasks and the saws have to be adjusted to perform cuts precisely according to the planned pattern. Every millimeter counts, when cutting such expensive material as ice cores.

At the same time, we have to monitor that each participating laboratory receives samples in the agreed amount and number.

What we have done today:

  1. Drilling with the NEEM long drill.
  2. Logging. Last bag: 483. Depth: 265.65 m.
  3. In the CFA lab. Peroxide and nitrate processes are now running.
  4. Cleaning up in camp. Removing pallets from main street.
  5. Building shelves in cooks freezer.
  6. Changing broken windows and doors in weatherports.
  7. Marking future position of storage garage.
  8. Working on a hardware problem on line scanner electronics.

Ad.1: This is how the drillers report looks today: A rather difficult day in the drill trench. First run with stable low current, 3 m long core and no packing in the pump. The second run was lost due to semi loose coupling and the third and last run was also aborted due to stuck pump between the fingers of the sleeve. It took some time to figure out what was going on. The borehole camera showed no damage to the sleeve. We need to mount thicker knobs to avoid further pump dislocations. The sharp edges of the brass booster have been filed down in order to avoid damage to the sleeve during mounting of the pump. The cutters are slightly worn and burred and were given polish on a very fine stone. The heads of the screws holding the 6 mm nominal pitch shoes were found to extrude slightly below the shoes. Thus the screws defined the pitch, not the shoes. Rufli filed down the heads. We will thus go back to the 4 mm shoes. We drilled 3.34 m today in 3 runs; drillers depth 252.90 m.

Weather: Fine all day with changing cloud cover. In the evening thin hazy overcast. -21°C to -10°C, 8-15 knots mainly from ESE. Visibility: Unrestricted.

Field leader, J.P. Steffensen

← Previous entry   Next entry →